OpenStudy (anonymous):

Solve x2 - 6x = 27 by completing the square. Show your work for full credit.

OpenStudy (cwrw238):

x2 - 6x = 27 (x - 3)^2 - 9 = 27 (x-3)^2 = 36 now take square root of both sides x - 3 = +/- 6 can you continue?

OpenStudy (anonymous):

Take half of the coefficient of the x, square it and add it to both sides. x^2-6x+9=27+9 (x+3)^2=36 take the square root x+3=sqrt(26) Separate into two problems x+3=+6 x+3=-6 x=6-3 x=-6-3 x=3 x=-9

OpenStudy (anonymous):

cwrw238 is correct it's (x-3)^2 not (x+3) my bad!

OpenStudy (cwrw238):

x - 3 = 6 x = 9 or x - 3 = --6 x = -3 solution set is {9,-3}

OpenStudy (anonymous):

Is that all? Please say yes